OpenStudy (anonymous):

simplify the expression sin^2theta/1+cos theta a. cos theta b.1+cos theta c.1-cos theta d.1-costheta/cos theta i think the answer is b

OpenStudy (neer2890):

sin^{2} θ =1−cos^ {2} θ=(1+cosθ)(1−cosθ)

OpenStudy (neer2890):

so c must be true

OpenStudy (anonymous):

\(\Large \frac{sin^{2}\theta}{1-cos\theta} = \frac{1-cos^{2}\theta}{1+cos\theta} = \frac{(1-cos\theta)(1+cos\theta)}{1+cos\theta} = 1-cos\theta \) Thus, answer is \(C\)